OpenStudy (anonymous):

How does lim as h goes to 0 (3+h) - s(3)/h = lim as h goes to 0 (3+h)^2 - 9/h? Instantaneous velocity for s(t)=t^2 at t=3.

OpenStudy (anonymous):

you don't want to plug in values until you've simplified it so you can take the limit without any problems...

OpenStudy (anonymous):

I don't understand how to simplify it :)

OpenStudy (anonymous):

f(t+h) = (t+h)^2 =t^2 +2th +h^2 -f(t) = -t^2

OpenStudy (anonymous):

( f(h+t) -f(t) ) / h = ( t^2 +2th +h^2 -t^2) / h

OpenStudy (anonymous):

you still can't take the limit as h->0 because h is in the denominator

OpenStudy (anonymous):

see any simplifications you can make in the numerator?

OpenStudy (anonymous):

factor an s?

OpenStudy (anonymous):

there are no s's there is a t^2 and a -t^2

OpenStudy (anonymous):

I just dont see it :/

OpenStudy (anonymous):

t^2 +2th +h^2 -t^2

OpenStudy (anonymous):

where? hold on maybe I didnt subtitute t^2 in?

OpenStudy (anonymous):

I don't know where to put the t^2 in the equation

OpenStudy (anonymous):

\[\frac{ s(t+h) -s(t) }{ h } = \frac{ t ^{2 } +2th +h^{2} -t ^{2}}{h }\]

OpenStudy (anonymous):

agree so far?

OpenStudy (anonymous):

hey what do I use to get the math formatted like that?

OpenStudy (anonymous):

equation editor

OpenStudy (anonymous):

thanks ok Im looking

OpenStudy (anonymous):

see the \[t ^{2} \] and the \[-t ^{2}\] in the numerator?

OpenStudy (anonymous):

Im not sure of the steps you took to get the equation on the right side of the =

OpenStudy (anonymous):

I plugged t+h into into s(t)

OpenStudy (anonymous):

subtracted s(t) from that

OpenStudy (anonymous):

and divided the whole thing by h

OpenStudy (anonymous):

s(t+h) = (t+h)^2 =t^2 +2th +h^2 -s(t) = -t^2

OpenStudy (anonymous):

ok Im screwing up on the pluggin in part :)

OpenStudy (anonymous):

Im confused by the t+h thing and trying to substitute

OpenStudy (anonymous):

put 't+h' in where ever there's a 't' in the equation this equation is simply 't^2' so it's easy: "(t+h)^2"

OpenStudy (anonymous):

yes Im seeing that now

OpenStudy (anonymous):

cool. you got it from here?

OpenStudy (anonymous):

and the 'S" just dissapears?

OpenStudy (anonymous):

because its not a number its like f(x)

OpenStudy (anonymous):

you're confused about function notation when they say s(t) = t^2 it's just a shorthand way of saying 's depends on t like t^2' ie if t=2 s=4 if t=9 s=81 and so forth

OpenStudy (anonymous):

the function is t^2.... that's what you want to deal with

OpenStudy (anonymous):

do you get t^2+3-t^2/3?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

that was just me substituting in t^2 and h=3

OpenStudy (anonymous):

\[\frac{ t^2+2th+h^2−t^2 }{ h }\]

OpenStudy (anonymous):

that's where we left off...

OpenStudy (anonymous):

need to simplify... t^2 and -t^2 cancel, so: \[\frac{2th+h^2 }{ h } \]

OpenStudy (anonymous):

the common term is 'h' \[ \frac{ h(2t+h) }{h }\]

OpenStudy (anonymous):

I dont understand how you get the first equation

OpenStudy (anonymous):

which?

OpenStudy (anonymous):

count 5 up from you

OpenStudy (anonymous):

I need to see from the substitution and then the VERY next step

OpenStudy (anonymous):

\[ \frac{t^2+2th+h^2−t^2 }{ h }\] that?

OpenStudy (anonymous):

yes im not substituting correctly

OpenStudy (anonymous):

I can't see what you're doing

OpenStudy (anonymous):

show me your steps or follow my steps.

OpenStudy (anonymous):

what does substituting t^2 into the equation look like?

OpenStudy (anonymous):

before ANY simplification :)

OpenStudy (anonymous):

naw man. you want to evaluate: \[\lim_{h \rightarrow 0} \frac{ s(t+h) -s(t) }{ h}\]

OpenStudy (anonymous):

yes ok thats the equation for instantaneous velocity

OpenStudy (anonymous):

find s(t+h) : put (t+h) into the equation wherever there's a 't' s(t) is simply t^2 so s(t+h) is (t+h)^2

OpenStudy (anonymous):

I get T^2 +h-t^2/h

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

\[(t+h)^2 = t^2 +2th +h^2\]

OpenStudy (anonymous):

for some reason its hard for me to see this

OpenStudy (anonymous):

foil it out and verify for yourself (t+h)(t+h) = ?

OpenStudy (anonymous):

(t+h)^2 makes sense Im used to seeing that but this? find s(t+h) : put (t+h) into the equation wherever there's a 't' s(t) is simply t^2 so s(t+h) is (t+h)^2

OpenStudy (anonymous):

that is killing me right now

OpenStudy (anonymous):

when you see s(4) what do you do?

OpenStudy (anonymous):

if s(t) = t^2

OpenStudy (anonymous):

you put 4 in for t

OpenStudy (anonymous):

s(4) =4^2 =16

OpenStudy (anonymous):

it's just notation to tell you what goes into the function s( whatever) = ( whatever)^2

OpenStudy (anonymous):

if we need s(t+h) then t+h goes in where ever there is a 't' and since s(t) is simply t^2 s(t+h) is simply (t+h)^2

OpenStudy (anonymous):

mull it over a bit

OpenStudy (anonymous):

internet just went out I mean power ewnt out